# A Stabilized Time-Domain Combined Field Integral Equation Using the Quasi-Helmholtz Projectors

### v1

@@ IV A stabilized TD-CFIE
We are in the position to propose a stabilized TD-CFIE by linearly combining the TD-EFIE (7) and the TD-MFIE (10). These two equations are compatible because they both map an RWG coefficient vector into the dual of the space spanned by the rotated BC functions. In particular, we consider the following combined formulation:
@@
Obviously, (11) is large-time step well-balanced, which scales as follows:
@@
Clearly, the choice of the coupling parameter has a major impact on the spectral properties of (11). In this paper, the optimal value $\eta^{2}$ is chosen, which minimizes the condition number of the system (11) for the electromagnetic scattering by the unit sphere [35]. Since the equation (11) is still time-dependent, a temporal discretization scheme is needed.
@@ IV-A Temporal discretization
Equation (11) is then tested with the Dirac delta distributions $\delta_{k}(t)=\delta(t-k\Delta t)$, with $k=1,2,\ldots,N_{T}$, resulting in the following lower triangular block matrix system:
@@ IV-B Marching-on-in-time algorithm
When the coefficient vectors $\operatorname{\mathbf{y}}_{i}$ of the auxiliary unknown $\mathrm{y}$ are computed, we can use the definition (5) and the expansion (12) to get back the physical current density $\mathrm{j}(t)$ as follows:
@@ V Numerical results
This section presents some numerical results for the scattering by PECs of two different geometries (see Fig. 5).
@@
The excitation is provided by the following Gaussian-in-time plane-wave electric field:
with amplitude $A=1\mathrm{V}$, width $w=26.67\mathrm{m}$, polarization $\boldsymbol{p}=\mathbf{1}_{x}$, direction $\boldsymbol{k}=\mathbf{1}_{z}$, and time of arrival $t_{0}=80\mathrm{ns}$. Accordingly, the incident magnetic field $\boldsymbol{h}^{in}$ is determined by
In order to examine the performance of the qHP TD-CFIE (11), its numerical results are compared with those of the following formulations:
@@
  • the mixed TD-CFIE proposed in [15], with $\alpha=0.5$.
These three equations are all temporally expanded in the set of hat functions $h_{i}(t)$, and then tested with the Dirac delta distributions $\delta_{k}(t)$. The coefficient vectors of the unknown $\boldsymbol{j}$ are computed using the time step $\Delta t=0.333\mathrm{ns}$.
@@ V-A Sphere
We start with the scattering by a simple PEC scatterer: a sphere of radius $1\mathrm{m}$ (see Fig. 5, left). It is approximated by 476 triangles with 240 vertices and 714 edges. The surface current density $\boldsymbol{j}(\boldsymbol{r},t)$ is computed using four different formulations, and its intensity at the point $\boldsymbol{r}=(-0.534,-0.523,-0.644)\mathrm{m}$ are presented in Fig. 6a. Before the resonances come into play (at $150\mathrm{ns}$, approximately), four equations yield similar results.
@@
Figure 7: Condition number of the matrix systems with respect to different mesh sizes: (a) of a sphere with radius $1\mathrm{m}$, while the time step is fixed at $\Delta t=3.33\mathrm{ns}$; (b) of a torus with two radii $3\mathrm{m}$ and $1\mathrm{m}$, while the time step is fixed at $\Delta t=26.667\mathrm{ns}$. The TD-EFIE and the mixed TD-CFIE become ill-conditioned when the mesh size decreases, whereas the TD-MFIE on the sphere and the proposed TD-CFIE stay well-conditioned. On toroidal surfaces, the TD-MFIE supports the non-trivial kernel of the static outer MFIE operator when the time step is sufficiently large.
@@ V-B Torus
The second experiment concerns the scattering by a PEC torus with large radius $3\mathrm{m}$ and small radius $1\mathrm{m}$ (see Fig. 5, right). The surface is discretized by 952 triangles with 1428 edges and 476 vertices. The intensity of the surface current density at the point $\boldsymbol{r}=(-2.326,-1.885,-0.692)\mathrm{m}$ are shown in Fig. 6b. Similar to the previous experiment, the TD-EFIE and the TD-MFIE are prone to resonant instabilities. Moreover, on toroidal surfaces, the TD-MFIE is also affected by near-dc instabilities rooted in the nullspaces of the static MFIE operators [6]. This effect can be noticed by the slow oscillation of the amplitude of the TD-MFIE solution. The TD-CFIEs are immune to resonant and dc instabilities. Therefore, their solutions decay, although slowly. From a very late time ($t>1000\mathrm{ns}$), the mixed TD-CFIE’s solution exhibits numerical noise, while the solution of the qHP TD-CFIE reveals a very slowly oscillating mode that takes place below the level of machine precision. This mode can be interpreted as the weak coupling of the TD-EFIE’s dc instabilities and the TD-MFIE’s near-dc instabilities. The coupling occurs because the harmonic components (global loops) exist and reside on the active range of both effects.
@@
Figures 7b and 8b again corroborate the well-conditioning of the proposed TD-CFIE (11) when the spatial mesh is fine or the time step is large. Please note that if $c\Delta t\geq D$ ($D$ is the diameter of the scatterer), the conditioning behavior of the TD-MFIE coincides with that of the static outer MFIE operator. This fact explains the increase of the condition number of the TD-MFIE along the resolution of the spatial mesh in Fig. 7b. It can also explain the behavior in Fig. 8b, where the condition number of the TD-MFIE firstly increases, and afterwards it stays constant when the time step is sufficiently large.

### v2

@@ IV A stabilized TD-CFIE
We are in the position to propose a stabilized TD-CFIE by linearly combining the TD-EFIE (7) and the TD-MFIE (10). These two equations are compatible because they both map an RWG coefficient vector into the dual of the space spanned by the rotated BC functions. In particular, we consider the following combined formulation
@@
Obviously, (11) is large-time step well-balanced, which scales as follows
@@
Clearly, the choice of the coupling parameter has a major impact on the spectral properties of (11). In this paper, the optimal value $\eta^{2}$ is chosen, which minimizes the condition number of the system (11) for the electromagnetic scattering by the unit sphere [42, 43]. Since the equation (11) is still time-dependent, a temporal discretization scheme is needed.
@@ IV-A Temporal discretization
Equation (11) is then tested with the Dirac delta distributions $\delta_{k}(t)=\delta(t-k\Delta t)$, with $k=1,2,\ldots,N_{T}$, resulting in the following lower triangular block matrix system
@@ IV-B Marching-on-in-time algorithm
When the coefficient vectors $\operatorname{\mathbf{y}}_{i}$ of the auxiliary unknown $\mathrm{y}$ are computed, we can use the definition (5) and the expansion (12) to get back the physical current density $\mathrm{j}(t)$ as follows
@@
IV-C Complexity analysis
The proposed qHP TD-CFIE involves rescaling and preconditioning procedures, necessitating additional matrix/vector product calculations compared to the standard TD-EFIE and TD-MFIE. Specifically, the rescaling involves the quasi-Helmholtz projectors, which can be constructed with linear complexity $\mathcal{O}(N_{E})$ [18]. Multiplication with these projectors also demands $\mathcal{O}(N_{E})$ operations. The preconditioning technique introduces the Gram matrix $\mathbb{G}$ and the Galerkin discretization matrices $\mathbb{Z}$ and $\mathbb{M}$ of the Yukawa-type operators $T_{-j\kappa}$ and $K_{-j\kappa}$. Whereas the former is sparse and well-conditioned, the latter are dense when the time step $\Delta t$ is large. Nevertheless, multiplication with them all have only a linear complexity $\mathcal{O}(N_{E})$, by means of acceleration techniques for quasi-static kernels [45]. Finally, the proposed MOT scheme can be coupled with fast time-domain algorithms, such as the plane-wave time-domain (PWTD) method [46], the time-domain adaptive integral method [47], the nonuniform grid time domain algorithm [48], and the accelerated Cartesian expansions [49], thereby reducing the computational cost even to scaling quasi-linearly with the number of spatial unknowns $N_{E}$ and linearly with the number of temporal unknowns $N_{T}$. For instance, the total complexity of the proposed numerical scheme is $\mathcal{O}(N_{T}N_{E}\log^{2}N_{E})$ using the PWTD method.
@@ V Numerical results
This section presents some numerical results for the scattering by PECs of two different geometries (see Fig. 5). The excitation is provided by a Gaussian-in-time plane-wave electric field of the form
with amplitude $A=1\mathrm{V}$, polarization $\boldsymbol{p}=\mathbf{1}_{x}$ and direction $\boldsymbol{k}=\mathbf{1}_{z}$. The width $w$ and the time of arrival $t_{0}$ are specified on each experiment. The incident magnetic field $\boldsymbol{h}^{in}$ is then determined by
@@
In order to examine the performance of the qHP TD-CFIE (11), its numerical results are compared with those of the following formulations
@@
  • the mixed TD-CFIE proposed in [15], with coupling parameter $\alpha=0.5$.
These three equations are all temporally expanded in the set of hat functions $h_{i}(t)$, and then tested with the Dirac delta distributions $\delta_{k}(t)$.
@@ V-A Sphere
We start with the scattering by a simple PEC scatterer: a sphere of radius $1\mathrm{m}$ (see Fig. 5, left). It is approximated by 476 triangles with 240 vertices and 714 edges.
Firstly, the exemption from the loss of accuracy at large time steps is verified. The Gaussian-in-time plane wave with $w=8\cdot 10^{6}\mathrm{m}$ and $t_{0}=80\mathrm{ms}$ is used as the excitation. The unknown surface current $\boldsymbol{j}(\boldsymbol{r},t)$ is numerically computed with time step $\Delta t=0.333\mathrm{ms}$, then it is Fourier transformed. Afterwards, the scattered far field computed from the transformed surface current at the frequency $3\mathrm{Hz}$ is compared with the reference one obtained by the Mie series. Fig. 6 shows that when the time step $\Delta t$ is large, the TD-EFIE’s and TD-MFIE’s solutions exhibit a loss of accuracy, resulting in inaccurate far fields. In contrast, the qHP TD-CFIE does not suffer from that loss.
Figure 6: Scattered far fields computed using the Mie series and different numerical schemes. Numerical solutions are computed with time step $\Delta t=0.333\mathrm{ms}$, then they are Fourier transformed at the frequency $3\mathrm{Hz}$. The TD-EFIE’s and TD-MFIE’s solutions are inaccurate due to a loss of significant digits on the irrotational components at large time steps, whereas the solution to the qHP TD-CFIE is accurate.
Next, we simulate the scattering by the PEC sphere using a moderate time step $\Delta t=0.333\mathrm{ns}$. Accordingly, the Gaussian pulse plane wave with width $w=8\mathrm{m}$ and time of arrival $t_{0}=80\mathrm{ns}$ is used. The surface current density $\boldsymbol{j}(\boldsymbol{r},t)$ is computed using four different formulations, and its intensity at the point $\boldsymbol{r}=(-0.534,-0.523,-0.644)\mathrm{m}$ are presented in Fig. 7a. Before the resonances come into play (at $150\mathrm{ns}$, approximately), four equations yield similar results.
@@
Figure 9: Condition number of the matrix systems with respect to different mesh sizes: (a) of a sphere with radius $1\mathrm{m}$, while the time step is fixed at $\Delta t=3.33\mathrm{ns}$; (b) of a torus with two radii $3\mathrm{m}$ and $1\mathrm{m}$, while the time step is fixed at $\Delta t=26.667\mathrm{ns}$. The TD-EFIE and the mixed TD-CFIE become ill-conditioned when the mesh size decreases, whereas the TD-MFIE on the sphere and the proposed qHP TD-CFIE stay well-conditioned. On toroidal surfaces, the TD-MFIE supports the non-trivial kernel of the static outer MFIE operator when the time step is sufficiently large.
@@ V-B Torus
The second experiment concerns the scattering by a PEC torus with large radius $3\mathrm{m}$ and small radius $1\mathrm{m}$ (see Fig. 5, right). The surface is discretized by 952 triangles with 1428 edges and 476 vertices. The incident plane wave is characterized by the width $w=8\mathrm{m}$ and the time of arrival $t_{0}=80\mathrm{ns}$, while the considered time step $\Delta t=0.333\mathrm{ns}$. The intensity of the surface current density at the point $\boldsymbol{r}=(-2.326,-1.885,-0.692)\mathrm{m}$ are shown in Fig. 7b. Similar to the previous experiment, the TD-EFIE and the TD-MFIE are prone to resonant instabilities. Moreover, on toroidal surfaces, the TD-MFIE is also affected by near-dc instabilities rooted in the nullspaces of the static MFIE operators [6]. This effect can be noticed by the slow oscillation of the amplitude of the TD-MFIE solution. The TD-CFIEs are immune to resonant and dc instabilities. Therefore, their solutions decay, although slowly. From a very late time ($t>1000\mathrm{ns}$), the mixed TD-CFIE’s solution exhibits numerical noise, while the solution of the qHP TD-CFIE reveals a very slowly oscillating mode that takes place below the level of machine precision. This mode can be interpreted as the weak coupling of the TD-EFIE’s dc instabilities and the TD-MFIE’s near-dc instabilities. The coupling occurs because the harmonic components (global loops) exist and reside on the active range of both effects.
@@
Figures 9b and 10b again corroborate the well-conditioning of the proposed TD-CFIE (11) when the spatial mesh is fine or the time step is large. Please note that if $c\Delta t\geq D$ ($D$ is the diameter of the scatterer), the conditioning behavior of the TD-MFIE coincides with that of the static outer MFIE operator. This fact explains the increase of the condition number of the TD-MFIE along the resolution of the spatial mesh in Fig. 9b. It can also explain the behavior in Fig. 10b, where the condition number of the TD-MFIE firstly increases, and afterwards it stays constant when the time step is sufficiently large.